OpenStudy (anonymous):

Divide. Express in lowest terms: 7/r divided by 63/r

OpenStudy (anonymous):

Div by 63/r is same as multiply by r/63....

OpenStudy (anonymous):

what estudier said it right. So to 'divide by 63/r' you can flip the fraction and multiply. 7/r * r/63 = 7/63 since the two r's cancel And then, you have 7/63 as a fraction which can be reduced down to 1/9 since 7*9 = 63